### v1

@@ Abstract
We tackle the hard problem of consciousness taking the naturally-selected, self-organising, embodied organism as our starting point. We provide a mathematical formalism describing how biological systems self-organise to hierarchically interpret unlabelled sensory information according to valence and specific needs. Such interpretations imply behavioural policies which can only be differentiated from each other by the qualitative aspect of information processing. Selection pressures favour systems that can intervene in the world to achieve homeostatic and reproductive goals. Quality is a property arising in such systems to link cause to affect to motivate real world interventions. This produces a range of qualitative classifiers (interoceptive and exteroceptive) that motivate specific actions and determine priorities and preferences. Building upon the seminal distinction between access and phenomenal consciousness, our radical claim here is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible. To put it provocatively: Nature does not like zombies. We formally describe the multilayered architecture of self-organisation from rocks to Einstein, illustrating how our argument applies in the real world. We claim that access consciousness at the human level is impossible without the ability to hierarchically model i) the self, ii) the world/others and iii) the self as modelled by others. Phenomenal consciousness is therefore required for human-level functionality. Our proposal lays the foundations of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to human fact than zombie fiction.
@@ 1 Introduction
Why is anything conscious? [^1] This work was supported by a Fundaçao para a Ciencia e a Tecnologia (FCT) grant PTDC/FER-FIL/4802/2020 ; and 2020-02773 CEECIND FCT to A.C. Both biological and other physical [^2] Physical here just means non-biological. We are not suggesting biological systems are non-physical. systems process information, yet it seems that humans consciously experience in addition to merely process information. Why? Living organisms are constantly processing self- and world-related information to secure survival in an ever-changing world. Human bodies share with all other physical systems the property of being instantiated in time and space, (e.g., our body occupies a given position and volume in space at a given time). Yet, unlike physical systems, biological, living systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [1, 2].
@@
Originally formalised in the field of cybernetics [3, 4] the notion of self-organisation has been subsequently applied to various disciplines including physics [5], biology [6, 7] and neuroscience [8, 9, 10]. Self-organization is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [11, 12, 13].
@@
Interestingly, much self- and world related information processing goes on behind the scenes, or “in the dark” so to speak, that is, without being constantly present to our conscious minds. But why doesn’t all information processing go in the dark?
@@
This question is the subject of long-standing debates across disciplines [14]. For example, one highly influential view is that consciousness has two aspects [14]. The first is functional, by which we mean the ability to access and communicate information [15]. How exactly information processing is linked to consciousness is the “easy problem” of consciousness [2]. The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [2, 15, 1, 16, 17]. This doesn’t mean just global states like being awake, but more specific local states like smelling a cup of coffee. These local contents or “qualia” are characterised by what it is like to be in them [14]. It is unclear the extent to which functional and phenomenal aspects are independent. David Chalmers has influentially suggested that it may be possible to construct a “zombie” which acts in every way like a person but has no qualia [15]. For example, a thermostat certainly detects heat and so processes information, but there presumably is not anything it is like to be a thermostat. Hence the question “why is anything conscious” may be understood as “why is there sometimes a qualitative aspect to information processing?”. This is the “hard problem” of consciousness.
@@
The hard problem has sparked a substantial body of work and a detailed discussion of these debates [14, 18] lies beyond the scope of our paper. Rather, in what follows we build upon the useful distinction between lower and higher order theories of information processing in relation to conscious experiences.
@@
For example, higher order thought theory (HOT) [19, 20] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states. Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those. The link between the two may explain something of the phenomenal character of states. Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness.
@@
The division between lower and higher orders is a good starting point to understand why some information processing goes on “in the dark”. Yet, the interesting question in our view is why do these lower order states arise? Can they occur in the absence of subjective, qualitative experience? It is widely agreed that higher order mental states such as desire to drink a coffee is accompanied by a qualitative aspect. But is there something it is like to “be in the dark”? That is: to process information at the lower bodily levels?
@@
In this paper we suggest that a useful way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
@@
If an explanation of how consciousness functions is to reveal why information processing sometimes has a qualitative aspect, then it must explain how we get local states [21], not begin by assuming local states. We need to go a level down and begin at the level of the embodied organism. Longtime considered a fringe approach, the embodied cognition paradigm [22] has recently gained substantial influence in cognitive science and philosophy [23, 24, 25]. The key idea is that instead of considering the body as a mere device designed to fuel and contain the mind (a device that can be replaced with a vat or a robot, for example), one must consider the mind as serving the self-sustaining needs of a surviving body.
@@
If this is so, then understanding consciousness must start with understanding the ‘humble’ lower bodily levels of information processing, and not the higher order levels of information processing only. The key idea is that conscious experiences do not merely depend on bodily experiences as an external factor that can be replaced with a vat or an artificial system.
@@
One promising approach to this question is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depend on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be functional [^3] Note that we are far from the first to claim that the phenomenal is functional. For example, the proposed Conscious Turing Machine [26] based on Global Workspace Theory [27], and constructivist approaches to artificial intelligence [28] take similar position. It is our explanation of how and why that is novel.. Unifying them. We would need to establish axioms that hold in every possible environment and show that it is impossible to have the function of consciousness without the subjective experience of it. Such an answer must explain how consciousness functions, and why some information processing goes on “in the dark”.
@@
This paper lays the grounds of a computational model formalizing the link between lower and higher levels of information processing in relation to conscious experiences in a way that is compatible with the basic principles of embodied cognition and enactivism. Enactivism is roughly the view that information processing arises through a dynamic interaction between an acting organism and its environment [29].
@@
Now, the notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [30, 31]. Here we define computation not in terms of symbol shuffling and representation, but in mechanistic terms. ‘Information processing’ in this sense is the causal relations dictating the transition of a system from one state to another, which concerns biological systems such as human bodies.
@@
Pancomputationalism is the idea that all dynamic systems are constituted by computation. This paper builds upon and extends the formalism of pancomputational enactivism developed by one of us [32]. To answer “why is anything conscious” we proceed in two steps. First we develop a mathematical formalism in which lower and higher order theories of consciousness, and phenomenal and access consciousness, are all derived from first principles. Each follows because of scaling natural selection pressures and the ability to adapt [^4] This is loosely inspired by a scale-based framing of machine learning [33].. Second, we provide an argument building on that to dissolve the hard problem of consciousness, answering why there must a qualitative aspect to information processing and in what circumstances.

### v2

@@ Abstract
We tackle the hard problem of consciousness taking the naturally selected, self-organising, embodied organism as our starting point. We provide a mathematical formalism describing how biological systems self-organise to hierarchically interpret unlabelled sensory information according to valence and specific needs. Such interpretations imply behavioural policies which can only be differentiated from each other by the qualitative aspect of information processing. Selection pressures favour systems that can intervene in the world to achieve homeostatic and reproductive goals. Quality is a property arising in such systems to link cause to affect to motivate real world interventions. This produces a range of qualitative classifiers (interoceptive and exteroceptive) that motivate specific actions and determine priorities and preferences. Building upon the seminal distinction between access and phenomenal consciousness, our radical claim here is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible. To put it provocatively: death grounds meaning, and Nature does not like zombies. We formally describe the multilayered architecture of self-organisation from rocks to Einstein, illustrating how our argument applies in the real world. We claim that access consciousness at the human level is impossible without the ability to hierarchically model i) the self, ii) the world/others and iii) the self as modelled by others. Phenomenal consciousness is therefore required for human-level functionality. Our proposal lays the foundations of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to human fact than zombie fiction.
@@ 1 Introduction
Why is anything conscious? Both biological and other physical [^1] Physical here just means non-biological. We are not suggesting biological systems are non-physical. systems process information, yet it seems that humans consciously experience in addition to merely processing information. Why? Living organisms are constantly processing self and world related information to survive in an ever-changing world. Human bodies share with all other physical systems the property of being instantiated in time and space, (e.g., our body occupies a given position and volume in space at a given time). Yet, unlike physical systems, biological, living systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [nagel1974, chalmers1995].
@@
Initially formally introduced in the field of cybernetics [ashby1947, foerster1960] the notion of self-organisation has been applied to various disciplines including physics [haken1983], biology [camazine2003, bell2007] and neuroscience [kelso1997, Friston2010, tognoli2014]. Self-organization is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [camazine2001, seeley2002, rosas2018].
@@
Interestingly, much self and world related information processing goes on behind the scenes, or “in the dark” so to speak, that is, without being constantly present to our conscious minds. But why doesn’t all information processing go in the dark?
@@
This question is the subject of long-standing debates across disciplines [seth2022]. One highly influential view is that consciousness has two aspects [seth2022]. The first is functional, by which we mean the ability to access and communicate information [block1995]. How information processing is linked to consciousness is the “easy problem” of consciousness [chalmers1995]. The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [chalmers1995, block1995, nagel1974, gallagher2021, fuchs2017]. This doesn’t mean just global states like being awake, but more specific local states like smelling a cup of coffee. These local contents or “qualia” are characterised by what it is like to be in them [seth2022]. It is unclear the extent to which functional and phenomenal aspects are independent. David Chalmers has influentially suggested that it may be possible to construct a “philosophical zombie” which acts in every way like a person but has no qualia [block1995]. For example, a thermostat certainly detects heat and so processes information, but there presumably is not anything it is like to be a thermostat. Hence the question “why is anything conscious” may be understood as “why is there sometimes a qualitative aspect to information processing?”. This is the “hard problem” of consciousness.
@@
The hard problem has sparked a substantial body of work, detailed discussion of which [seth2022, northoff2014] lies beyond the scope of our paper. Rather, in what follows we build upon the useful distinction between lower and higher order theories of information processing in relation to conscious experiences.
@@
For example, higher order thought theory (HOT) [rosenthal2005, brown2019] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states. Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those. The link between the two may explain something of the phenomenal character of states. Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness.
@@
The distinction between lower and higher orders is a good starting point to understand why most biological information processing goes on “in the dark”. Yet, the interesting question in our view is why do these lower order states arise? Can they occur in the absence of subjective, qualitative experience? It is widely agreed that higher order mental states such as desire to drink a coffee is accompanied by a qualitative aspect. But is there something it is like to “be in the dark”? That is: to process information at the lower bodily levels?
@@
In this paper we suggest a way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
@@
If an explanation of how consciousness functions is to reveal why information processing sometimes has a qualitative aspect, then it must explain how we get local states [thompson2016], not begin by assuming local states. We need to go a level down and begin at the level of the embodied organism. Longtime considered a fringe approach, the embodied cognition paradigm [varela2016] has recently gained substantial influence in cognitive science and philosophy [ciaunica2023, levin2024, seth2018]. The idea is that instead of considering the body as a mere device designed to fuel and contain the mind (a device that can be replaced with a vat or a robot, for example), one must consider the mind as serving the self-sustaining needs of a surviving body.
@@
If this is so then understanding consciousness must start with understanding the ‘humble’ lower bodily levels of information processing, not the higher order levels of information processing only. In other words, conscious experiences do not merely depend on bodily experiences as an external factor that can be replaced with a vat or an artificial system.
@@
One promising approach to this question is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depend on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be intrinsically functional [^2] Note that we are far from the first to claim that the phenomenal is functional. For example, the proposed Conscious Turing Machine [blum2020] based on Global Workspace Theory [baars1997], and constructivist approaches to artificial intelligence [wang2020] take similar position. It is our explanation of how and why that is novel.. Unifying them. We would need to establish axioms that hold in every conceivable environment and show that it is impossible to have the function of consciousness without the subjective experience of it. Such an answer must explain how consciousness functions, and why most information processing goes on “in the dark”.
@@
This paper lays the grounds of a computational model formalizing the link between lower and higher levels of information processing in relation to conscious experiences in a way that is compatible with the basic principles of embodied cognition and enactivism. Enactivism is roughly the view that information processing arises through a dynamic interaction between an acting organism and its environment [thompson2007].
@@
Now, the notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [piccinini2015, piccinini2021]. Here we define computation not in terms of abstract symbol shuffling and representation, but in concrete mechanistic terms where ‘information processing’ refers to the actual physical transition of a system from one state to another; a system such as a human body.
@@
Pancomputationalism is the idea that all dynamic systems are constituted by computation, which for our purposes is defined as above. This paper builds upon the formalism of pancomputational enactivism [bennett2024a]. To answer “why is anything conscious” we proceed in two steps. First, we develop a formalism in which lower and higher order theories of consciousness, and phenomenal and access consciousness, are all derived from first principles. Each follows because of scaling natural selection pressures and the ability to adapt [^3] This is loosely inspired by a scale-based framing of machine learning [sutton2019] - we are applying Sutton’s “bitter lesson” to biology.. Second, we provide an argument building on that to dissolve the hard problem of consciousness, answering why there must a qualitative aspect to information processing and in what circumstances.